I am having a very difficult time with part of function notation problems.
f(x)=-2x + 4
find f(-1/5)
I need first grade example of how the 4 ends up as 20/5.

Found 2 solutions by mananth, Alan3354:
Answer by mananth(16949)

f(x)=-2x + 4
find f(-1/5)
f(-1/5) = 2(-1/5)+4
f(-1/5)= (-2/5) +4
the LCD = 5

((-2*5 + 5*4)/5
=-2/5 +20/5
=18/5

Answer by Alan3354(69443)

4*(5/5) = 20/5
You can always multiply or divide by 1.
1 = 5/5, or -6/-6, or n/n if n is not = 0.